# GW231123: a Binary Black Hole Merger with Total Mass 190-265

### v2

@@ 6.1 Single-event rate estimate
First, we quantify the merger rate of GW231123-like events following (Kim et al., 2003; Abbott et al., 2016b). The sensitive volume–time of the detectors to such signals is estimated using the cWB-BBH results for the injection campaign discussed in Section 2. Assuming a constant merger rate over comoving volume and source-frame time with prior and a Poisson likelihood for the number of triggers, we find . This is consistent with the rate of mergers like GW190521 and upper limits of IMBH mergers (Abbott et al., 2020d, e, 2022), but much lower than the overall rate 16–61 Gpc-3 yr-1 of BBHs with component masses inferred through GWTC-3 (Abbott et al., 2023b).
@@ 6.2 Relation to the previous inferred population
To further assess GW231123 in the context of the 69 BBH mergers with through GWTC-3 (Abbott et al., 2023a) and test if its masses and spins are surprising, we perform posterior predictive checks based on the fiducial BBH population fit from (Abbott et al., 2023b, Section III C; we extend the prior on the maximum BH mass up to as there is support from GW190521 above the limit of 100 imposed in the original GWTC-3 analysis). From the inferred population, we construct mock catalogs containing 69 detected events and plot the distribution of their largest BH mass in Figure 3. Using the Combined parameter estimates, the primary mass of GW231123 falls at the level of this distribution, indeed indicating that this event is an unlikely draw. However, due to large uncertainties in its masses, it is not conclusively an outlier as it may be less massive than the most massive mock events (equivalent comparisons for secondary and total mass are less significant). Similarly, the largest BH spin in these catalogs is , against which the primary and secondary spins of GW231123 can fall at any percentile and thus are not outliers. Compared to its masses, the spins of GW231123’s source are more consistent with the known population as it does not rule out large values.
@@ 6.3 Possible formation channels
From theoretical predictions for the late-stage evolution of massive stars (Fowler & Hoyle, 1964; Barkat et al., 1967; Rakavy & Shaviv, 1967; Fraley, 1968; Bond et al., 1984; Woosley et al., 2002; Woosley, 2017), contraction of the core leads to electron–positron pair production that reduces internal pressure support, causing further contraction that powers explosive nuclear burning and a rebounding shock. For helium-core masses –, multiple pulsational episodes can eject sufficient material to reduce the mass below the pair-instability regime, ending with a BH remnant. A single pulse can entirely disrupt stars with larger helium cores, leaving behind no remnant in a PISN. At even larger helium-core masses , this is avoided as the high core temperature results in photodisintegration that accelerates gravitational collapse to a massive BH. This leads to the robust prediction from single-star evolution of the existence of a gap in the BH mass distribution. Though this gap is broadly consistent with the range –, there are several theoretical uncertainties that affect both the lower edge and total extent of the gap (Belczynski et al., 2016; Stevenson et al., 2019; Farmer et al., 2019; Mapelli et al., 2020; Renzo et al., 2020b; Marchant & Moriya, 2020; Woosley & Heger, 2021; Hendriks et al., 2023). Uncertainties in nuclear reaction rates alone can shift the lower edge of the pair-instability mass gap from to (Farmer et al., 2020).
@@
Some stellar and binary evolution processes are predicted to be able to populate the pair-instability mass gap. Weaker stellar winds (Mapelli et al., 2020) or core dredge-up episodes (Costa et al., 2021) may allow a star to retain a hydrogen envelope and collapse to a BH with mass inside the gap (Spera et al., 2019). Stellar mergers in dense clusters or multiple systems could similarly produce stars that collapse to more massive BHs, with the large masses inferred from GW231123 perhaps requiring multiple such mergers (Mapelli, 2016; Di Carlo et al., 2020; Renzo et al., 2020a; Kremer et al., 2020; González et al., 2021; Rizzuto et al., 2022; Costa et al., 2022; Arca Sedda et al., 2023a). Short-period stellar binaries might avoid merging and produce binary BHs with large, equal masses from rapidly rotating metal-poor stars due to chemically homogeneous evolution (de Mink & Mandel, 2016; Mandel & de Mink, 2016; Marchant et al., 2016). However, most models of isolated-binary formation predict small natal spins and at most one of the BHs spinning, due to tidal synchronization or accretion-induced spin up, and so binaries with masses and spins like those inferred from GW231123 are difficult to form (Belczynski et al., 2020; Qin et al., 2018; Fuller & Ma, 2019; Bavera et al., 2020; Belczynski et al., 2020; van Son et al., 2020). In fact, the components BHs and especially the primary are so massive that they may have formed through core collapse above the pair-instability mass gap (Ezquiaga & Holz, 2021; Franciolini et al., 2024).
@@
Alternatively, in hierarchical mergers, one or both of the binary components is the product of a previous BBH merger, with characteristically large masses and spins (Gerosa & Fishbach, 2021). Previous analyses have suggested evidence for hierarchical mergers in GW catalogs (Kimball et al., 2021; Mould et al., 2022; Wang et al., 2022; Li et al., 2024; Pierra et al., 2024; Hussain et al., 2024; Antonini et al., 2025), but the population of BH remnants receive gravitational recoils as high as – (Doctor et al., 2021; Mahapatra et al., 2021), requiring environments with high escape speeds (Antonini & Rasio, 2016) such as dense stellar clusters (Miller & Hamilton, 2002; Antonini et al., 2019; Rodriguez et al., 2019; Fragione & Silk, 2020; Mapelli et al., 2021; Arca Sedda et al., 2021; Kritos et al., 2023; Mahapatra et al., 2025) or the disks of active galactic nuclei (Bartos et al., 2017; Stone et al., 2017; Mckernan et al., 2018; Yang et al., 2019; Tagawa et al., 2020; McKernan et al., 2020; Vaccaro et al., 2024; Arca Sedda et al., 2023b) to be retained. This is in contrast to high-mass stellar mergers, which receive smaller recoils from asymmetric mass loss (Gaburov et al., 2010; Glebbeek et al., 2013) and therefore may be more efficient at producing BHs with large masses in dynamical environments. As seen in Figure 4, the spins inferred from GW231123 may be even larger than typically predicted from hierarchical BH mergers (Gerosa & Berti, 2017; Fishbach et al., 2017), although the expected distribution for sources retained in their host environments may accommodate a wider range of spins (Borchers et al., 2025).
@@
A different possibility is that of primordial BH being the binary components, which may exist across a range of mass scales, including within the pair-instability mass gap (Bird et al., 2016, 2023; Clesse & Garc´ıa-Bellido, 2017, 2022). However, there are remaining theoretical uncertainties, e.g., on whether primordial BH could accrete sufficiently to spin up as rapidly as the BH inferred from GW231123 (Green & Kavanagh, 2021).
@@
Altogether, these theoretical predictions and their uncertainties make it difficult to determine whether or not the BH in the source of GW231123 have an astrophysical origin directly from stellar collapse, or due to close encounters in dense clusters (Jaraba & Garc´ıa-Bellido, 2021). We quantify this in more detail below.
@@ 6.4 Stellar collapse
To account for a range of possible locations for the pair-instability mass gap, in Figure 6 we compute the probability that one or both component masses fall within the gap as a function of its lower edge from 40–100 and upper edge from 120–180 , using the Combined parameter estimates. In the following, we quote these probabilities specifically for the putative gap 60–130 . The probabilities that the secondary (primary) BH lies in, above, and below this gap are 83 % (26 %), 1 % (74 %), and 16 % (0 %), respectively. Considering scenarios in which at least one of the components falls in this gap, the joint probability that: both BH are in the gap (upper left panel of Figure 6) is 25 %; the primary is above while the secondary is within (upper right) is 58 %; and that the primary is within while the secondary is below (lower left) is 1 %. Alternatively, a scenario with neither BH in the gap is possible in the case of a straddling binary (Fishbach & Holz, 2020), with a primary BH above the upper edge and secondary below the lower edge (lower right) having a probability of 14 %.
@@ 6.5 Hierarchical mergers
Given the high probability of at least one of the BH lying inside the pair-instability mass gap, we consider the possibility that this is due to repeated BBH mergers. Assuming hierarchical origins, several works have inferred the source properties of potential BBHs whose merger products are observed with GWs in a subsequent merger (Baibhav et al., 2021; Barrera & Bartos, 2022; Álvarez et al., 2024; Mahapatra et al., 2024). We follow (Álvarez et al., 2024) and use the NRSur7dq4Remnant surrogate model (Varma et al., 2019) to find the distribution of BBH source properties such that the corresponding distribution of BH remnant properties reproduces the Combined posterior over mass and spin for the primary and secondary BH inferred from GW231123. As the primary-spin posterior favors large values , this constrains the parent binary of the primary BH to have unequal masses and ; for more equal masses, both BH spins can reduce the total angular momentum if misaligned, whereas unequal-mass binaries are dominated by the single heavier BH. The parent binary of the primary BH may have had a large effective inspiral spin, with , but is not ruled out. A similar picture holds for the secondary BH, with parent masses and , but more uncertain effective inspiral spin due to the larger uncertainty on the secondary spin in the source of GW231123. These mergers would have imparted kicks of and in the case of the primary and secondary, respectively, resulting in ejection from young star clusters or globular clusters (Antonini & Rasio, 2016).
The heavier of the two BH in both parent binaries may also lie within the pair-instability mass gap, with probabilities 91 % and 74 % for the heavier parent of the primary and secondary, respectively, when taking a gap from 60–130 , as above. Therefore, if either of the component BH of GW231123’s source is interpreted as the product of a previous BH merger, it may be the result of multiple previous mergers. Alternatively, the parent BH could themselves have formed with large masses by the other astrophysical processes discussed above, such as stellar mergers.
@@ 7 Alternative interpretations
All GW observations to date have been inferred to be from compact binaries consisting of BH and/or neutron stars (Abbott et al., 2016c, 2021a, 2024a, 2023a; Venumadhav et al., 2020; Olsen et al., 2022; Mehta et al., 2025; Nitz et al., 2023; Wadekar et al., 2023), and we consider a BBH the most astrophysically plausible interpretation of GW231123, finding that a non-eccentric BBH model fits the signal with no significant residual. Nonetheless, the low number of observable GW cycles invites alternative interpretations. We discuss several here.
@@ 7.1 Eccentricity
Binaries formed in dense environments may retain residual eccentricity in the sensitive band of current GW detectors (Samsing, 2018; Rodriguez et al., 2018; Zevin et al., 2019; Chattopadhyay et al., 2023; Dall’Amico et al., 2024) or merge following a dynamical capture (Gold & Brügmann, 2013; East et al., 2013; Gamba et al., 2023; Andrade et al., 2024; Albanesi et al., 2025b), but for high masses their GW signals can be confused with those of non-eccentric mergers (Romero-Shaw et al., 2020a; Calderón Bustillo et al., 2021a; Romero-Shaw et al., 2023). Our signal models assume a non-eccentric inspiral, while state-of-the-art IMR models that include eccentricity (Liu et al., 2022; Gamboa et al., 2024; Paul et al., 2025; Albanesi et al., 2025a; Planas et al., 2025a) assume circularization in the merger–ringdown stages and would thus be unsuitable to infer the parameters of GW231123’s source if it was eccentric when observed (Ramos-Buades et al., 2023b; Iglesias et al., 2024; Gupte et al., 2024; Planas et al., 2025b). Extensions of QNM amplitude models beyond eccentric non-spinning configurations (Carullo, 2024) will be required to investigate the possible ringdown mode excitation hinted at in Section 5. Relaxing the non-eccentric assumption is not expected to significantly change our results unless the eccentricity is larger than close to merger (Hinder et al., 2008; Huerta et al., 2019; Healy & Lousto, 2022; Carullo et al., 2024; Nee et al., 2025), which would be rare in the dynamical-capture scenarios above. Therefore, although we consider large eccentricity at merger astrophysically unlikely, we cannot rule it out for the source of GW231123.
@@ 7.2 Gravitational lensing
GW signals may be strongly lensed by galaxies or galaxy clusters, producing multiple copies of the original signal (Hannuksela et al., 2019; Abbott et al., 2021c, 2024b). However, no closely matching super-threshold counterpart candidates for GW231123 have been found from standard CBC searches. GWs can also undergo wave-optics lensing (Takahashi & Nakamura, 2003) when they encounter smaller objects (– for signals in the LVK band). GW231123 shows the strongest support for distorted lensed signals seen so far for both a point-mass model (Wright & Hendry, 2022) and phenomenological analyses (Liu et al., 2023), although preliminary background analyses suggest that some GW231123-like signals may be mis-identified as lensed. More in-depth investigations are needed to assess the significance of the lensing hypothesis, and these will be presented in future work.

### v3

@@ 6.1 Single-event rate estimate
First, we quantify the merger rate of GW231123-like events following (Kim:2002uw; LIGOScientific:2016kwr). The sensitive volume–time of the detectors to signals whose source properties are consistent with the posterior distribution of GW231123 is estimated using the cWB-BBH results for the injection campaign in Section 2. Assuming a constant merger rate over comoving volume and source-frame time with prior and a Poisson likelihood for the number of triggers, we find . This is consistent with the rate of mergers like GW190521 (; LIGOScientific:2020iuh; LIGOScientific:2020ufj) and upper limits of IMBH mergers (e.g., even the most stringent 90% upper limit from LIGOScientific:2021tfm and see Table 3 therein for constraints across source properties), but lower than the overall rate of BBHs with component masses inferred through GWTC-3 (16–61 Gpc-3 yr-1; KAGRA:2021duu).
@@ 6.2 Relation to the previous inferred population
To further assess GW231123 in the context of the 69 BBH mergers with through GWTC-3 (LIGOScientific:2021vkt) and test if its masses and spins are surprising, we perform posterior predictive checks based on the fiducial BBH population fit from (KAGRA:2021duu, Section III C; we extend the prior on the maximum BH mass up to as there is support from GW190521 above the limit of 100 imposed in the original GWTC-3 analysis). From the inferred population, we construct mock catalogs containing 69 detected events and plot the distribution of their largest BH mass in Figure 3. Using the combined parameter estimates in Section 4.3, the primary mass of GW231123 falls at the level of this distribution, indeed indicating that this event is an unlikely draw. However, due to large uncertainties in its masses, it is not conclusively an outlier as it may be less massive than the most massive mock events (equivalent comparisons for secondary and total mass are less significant). Similarly, the largest BH spin in these catalogs is , against which the primary and secondary spins of GW231123 can fall at any percentile and thus are not outliers. Compared to its masses, the spins of GW231123’s source are more consistent with the known population as it does not rule out large values.
@@ 6.3 Possible formation channels
From theoretical predictions for the late-stage evolution of massive stars (Fowler:1964zz; Barkat:1967zz; 1967ApJ...148..803R; 1968Ap&SS...2...96F; Bond:1984sn; Woosley:2002zz; Woosley:2016hmi), contraction of the core leads to electron–positron pair production that reduces internal pressure support, causing further contraction that powers explosive nuclear burning and a rebounding shock. For helium-core masses –, multiple pulsational episodes can eject sufficient material to reduce the mass below the pair-instability regime, ending with a BH remnant. A single pulse can entirely disrupt stars with larger helium cores, leaving behind no remnant in a PISN. At even larger helium-core masses , this is avoided as the high core temperature results in photodisintegration that accelerates gravitational collapse to a massive BH. This leads to the robust prediction from single-star evolution of the existence of a gap in the BH mass distribution. Though this gap is broadly consistent with the range –, there are several theoretical uncertainties that affect both the lower edge and total extent of the gap (Belczynski:2016jno; Stevenson:2019rcw; 2019ApJ...887...53F; Mapelli:2019ipt; Renzo:2020rzx; Marchant:2020haw; Woosley:2021xba; Hendriks:2023yrw). Uncertainties in nuclear reaction rates alone can shift the lower edge of the pair-instability mass gap from to (Farmer:2020xne).
@@
Some stellar and binary evolution processes are predicted to be able to populate the pair-instability mass gap. Weaker stellar winds (Mapelli:2019ipt) or core dredge-up episodes (Costa:2020xbc) may allow a star to retain a hydrogen envelope and collapse to a BH with mass inside the gap (Spera:2018wnw). Short-period stellar binaries might avoid merging and produce binary BHs with large, equal masses from rapidly rotating metal-poor stars due to chemically homogeneous evolution (deMink:2016vkw; Mandel:2015qlu; Marchant:2016wow). However, most models of isolated-binary formation predict small natal spins and at most one of the BHs spinning, due to tidal synchronization or accretion-induced spin up, and so binaries with masses and spins like those inferred from GW231123 are difficult to form (Belczynski:2017gds; Qin:2018vaa; Fuller:2019sxi; Bavera:2020inc; Belczynski:2017gds; vanSon:2020zbk). In fact, the components BHs and especially the primary are so massive that they may have formed through core collapse above the pair-instability mass gap (Ezquiaga:2020tns; Franciolini:2024vis).
@@
Alternatively, in hierarchical mergers, one or both of the binary components is the product of a previous BBH merger, with characteristically large masses and spins (Gerosa:2021mno). As seen in Figure 4, the spins inferred from GW231123 may be even larger than typically predicted from hierarchical BH mergers (Gerosa:2017kvu; Fishbach:2017dwv), although the expected distribution for sources retained in their host environments may accommodate a wider range of spins (Borchers:2025sid). Previous analyses have suggested evidence for hierarchical mergers in GW catalogs (Kimball:2020qyd; Mould:2022ccw; Wang:2022gnx; Li:2023yyt; Pierra:2024fbl; Hussain:2024qzl; Antonini:2024het), but the population of BH remnants receive gravitational recoils as high as – (Doctor:2021qfn; Mahapatra:2021hme), requiring environments with high escape speeds (Antonini:2016gqe) such as dense stellar clusters (Miller:2001ez; Antonini:2018auk; Rodriguez:2019huv; Fragione:2020nib; Mapelli:2020xeq; Sedda:2021abh; Kritos:2022non; Mahapatra:2022ngs) or active galactic nucleus (AGN) disks (Bartos:2016dgn; Stone:2016wzz; Mckernan:2017ssq; Yang:2019cbr; Tagawa:2019osr; McKernan:2019beu; Vaccaro:2023cwr; Sedda:2023big) to be retained.
This is in contrast to stellar mergers, which receive smaller recoils from asymmetric mass loss (Gaburov:2009kg; Glebbeek:2013kga) and therefore may be more efficient at producing BH with large masses in dynamical environments. The large masses inferred from GW231123 may be explained by multiple such mergers in dense clusters or multiple systems (Mapelli:2016vca; DiCarlo:2019fcq; Renzo:2020smh; Kremer:2020wtp; Gonzalez:2020xah; Rizzuto:2021atw; Costa:2022aka; ArcaSedda:2023mlv), a scenario that could also describe the formation of massive central BH (PortegiesZwart:2004ggg; Greene:2019vlv).
Besides mass transfer between the components of a stellar binary, BH mass growth may also occur via accretion in other gaseous environments. BH embedded in the disk of an AGN may accrete material directly from the disk or from collisions with disk stars (McKernan:2012rf). Similarly, in dense clusters, BH may accrete from stars after undoing dynamical interactions (Giersz:2015mlk; Lopez:2018nkj; Kiroglu:2024xpc). Furthermore, these accretion processes may also increase BH spins.
@@
A different possibility is that of primordial BH being the binary components, which may exist across a range of mass scales, including within the pair-instability mass gap (Bird:2016dcv; Bird:2022wvk; Clesse:2016vqa; Clesse:2020ghq). However, there are remaining theoretical uncertainties, e.g., on whether primordial BH could accrete sufficiently to spin up as rapidly as the BH inferred from GW231123 (Green:2020jor).
@@
Altogether, these theoretical predictions and their uncertainties make it difficult to determine whether or not the BH in the source of GW231123 have an astrophysical origin directly from stellar collapse. We quantify this in more detail below.
@@ 6.4 Stellar collapse
To account for a range of possible locations for the pair-instability mass gap, in Figure 6 we compute the probability that one or both component masses fall within the gap as a function of its lower edge from 40–100 and upper edge from 120–180 , using the combined parameter estimates. In the following, we quote these probabilities specifically for the putative gap 60–130 . The probabilities that the secondary (primary) BH lies in, above, and below this gap are 83 % (28 %), 1 % (72 %), and 16 % (0 %), respectively. Considering scenarios in which at least one of the components falls in this gap, the joint probability that: both BH are in the gap (upper left panel of Figure 6) is 26 %; the primary is above while the secondary is within (upper right) is 57 %; and that the primary is within while the secondary is below (lower left) is 2 %. Alternatively, a scenario with neither BH in the gap is possible in the case of a straddling binary (Fishbach:2020qag), with a primary BH above the upper edge and secondary below the lower edge (lower right) having a probability of 14 %.
@@ 6.5 Hierarchical mergers
Given the high probability of at least one of the BH lying inside the pair-instability mass gap, we consider the possibility that this is due to repeated BBH mergers. Assuming hierarchical origins, several works have inferred the source properties of potential BBHs whose merger products are observed with GWs in a subsequent merger (Baibhav:2021qzw; Barrera:2022yfj; Alvarez:2024dpd; Mahapatra:2024qsy). We follow (Alvarez:2024dpd) and use the NRSur7dq4Remnant surrogate model (Varma:2019csw) to find the distribution of BBH source properties such that the corresponding distribution of BH remnant properties reproduces the combined parameter estimates over mass and spin for the primary and secondary BH inferred from GW231123. As the primary-spin posterior favors large values , this constrains the parent binary of the primary BH to have unequal masses and ; for more equal masses, both BH spins can reduce the total angular momentum if misaligned, whereas unequal-mass binaries are dominated by the single heavier BH. The parent binary of the primary BH may have had a large effective inspiral spin, with , but is not ruled out. A similar picture holds for the secondary BH, with parent masses and , but more uncertain effective inspiral spin due to the larger uncertainty on the secondary spin in the source of GW231123. These mergers would have imparted kicks of and in the case of the primary and secondary, respectively, resulting in ejection from environments with escape speeds , such as young star clusters or globular clusters (Antonini:2016gqe).
The heavier of the two BH in both parent binaries may also lie within the pair-instability mass gap, with probabilities 96 % and 71 % for the heavier parent of the primary and secondary, respectively, when taking a gap from 60–130 , as above. Therefore, if either of the component BH of GW231123’s source is interpreted as the product of a previous BH merger, it may be the result of multiple previous mergers or require the components of the parent binary to have formed with larger masses via other astrophysical processes, such as stellar mergers or BH accretion, as discussed in Section 6.3.
@@ 7 Alternative interpretations
All GW observations to date have been inferred to be from compact binaries consisting of BH and/or neutron stars (LIGOScientific:2016dsl; LIGOScientific:2020ibl; LIGOScientific:2021usb; LIGOScientific:2021vkt; Venumadhav:2019lyq; Olsen:2022pin; Mehta:2023zlk; Nitz:2021zwj; Wadekar:2023gea), and we consider a BBH the most astrophysically plausible interpretation of GW231123, finding that a non-eccentric BBH model fits the signal with no significant residual. Nonetheless, the low number of observable GW cycles invites alternative interpretations. We discuss several here.
@@ 7.1 Eccentricity
Binaries formed in dense environments may retain residual eccentricity in the sensitive band of current GW detectors (Antonini:2013tea; Samsing:2017xmd; Rodriguez:2017pec; Zevin:2018kzq; Chattopadhyay:2023pil; DallAmico:2023neb) or form with large eccentricities and merge promptly after due to a dynamical capture (Gold:2012tk; East:2012xq; Gamba:2021gap; Andrade:2023trh; Albanesi:2024xus), but for high masses their GW signals can be confused with those of non-eccentric mergers (Romero-Shaw:2020thy; CalderonBustillo:2020xms; Romero-Shaw:2022fbf). Our signal models assume a non-eccentric inspiral, while state-of-the-art IMR models that include eccentricity (Liu:2021pkr; Gamboa:2024hli; Paul:2024ujx; Albanesi:2025txj; Planas:2025feq) assume circularization in the merger–ringdown stages and would thus be unsuitable to infer the parameters of GW231123’s source if it was eccentric when observed (Ramos-Buades:2023yhy; Iglesias:2022xfc; Gupte:2024jfe; Planas:2025jny). Extensions of QNM amplitude models beyond eccentric non-spinning configurations (Carullo:2024smg) will be required to investigate the possible ringdown mode excitation hinted at in Section 5. Many studies have found that the merger-ringdown signal is robust with respect to moderate inspiral eccentricity (Hinder:2007qu; Huerta:2019oxn; Healy:2022wdn; Carullo:2023kvj; Nee:2025zdy). Relaxing the non-eccentric assumption is not expected to significantly change our results unless the eccentricity is larger than close to merger (Healy:2022wdn; Carullo:2023kvj), which would be rare in the dynamical-capture scenarios above. For example, Chattopadhyay:2023pil find an overall merger rate in dense stellar clusters, 10% with eccentricity at a GW frequency of 10 Hz and 10% involving BH with masses , implying a rate of massive eccentric mergers , already at the lower limit of our constraint for GW231123 without considering the decline in the number of sources at increasing mass and eccentricity. Although we do not explicitly rule out large eccentricity for the source of GW231123, we therefore consider it astrophysically unlikely.
@@ 7.2 Gravitational lensing
GW signals may be strongly lensed by galaxies or galaxy clusters, producing multiple copies of the original signal (Hannuksela:2019kle; LIGOScientific:2021izm; LIGOScientific:2023bwz). However, no closely matching super-threshold counterpart candidates for GW231123 have been found from standard CBC searches. GWs can also undergo wave-optics lensing (Takahashi:2003ix) when they encounter smaller objects (– for signals in the LVK band). GW231123 shows the strongest support for distorted lensed signals seen so far for both a point-mass model (Wright_2022) and phenomenological analyses (Liu:2023ikc), although preliminary background analyses suggest that some GW231123-like signals may be mis-identified as lensed. More in-depth investigations are needed to assess the significance of the lensing hypothesis, and these will be presented in future work.